# The Case for ‘Thick Evaluations’ of Cultural Representation in AI

### v1

@@ 3. Methods
In this research, we treat the evaluation process as an object of study itself. We engaged with participants from three South Asian countries (Sri Lanka, Pakistan, and India) through a series of individual online evaluations and collective in-person workshops. These methods were chosen to allow us to move beyond simplistic evaluation metrics; we instead sought to capture in-depth, qualitative reflections and discussions participants have as they interpret and evaluate AI-generated images. Given the discursive and collective nature of representation, we chose group workshops instead of individual interviews. To preserve participants’ individual perspectives, we included an individual evaluation task prior to the workshops. Additionally, to ensure the annotation process reflected local cultural norms—including platform choice, framing of questions, and participant composition—we collaborated with local research partners in each region.
3.1. Participants and Sites
@@
We chose our research sites for two main reasons. First, we aim to contribute to a growing body of scholarship that seeks to expand the evaluation of AI beyond Western-centric perspectives (Kak, 2020; Mohamed et al., 2020; Sambasivan et al., 2021; Qadri et al., 2023a). Second, we intend to move beyond the common “USA vs. India” dichotomy often found in comparative analyses including the Global South (e.g., (Raman et al., 2008; Khairullah and Khairullah, 2009)). To this end, we focused on three South Asian countries: Sri Lanka, Pakistan, and India. They were selected due to the shared cultural histories, while also recognizing the vast diversity within South Asia. We recruited participants through collaborations with local research partners deeply embedded in their communities. Our partners in Sri Lanka and Pakistan were university researchers with extensive research experience with marginalized local communities, while our partner in India was a research manager experienced in conducting workshops and focus groups. Recruitment began with emails to targeted lists, including local institutions, previous research participants that worked with those partners, and the partners’ professional networks. Interested individuals completed a screening form detailing demographics and background.
@@
Rather than attempting to exhaustively represent any single culture, our aim was to capture nuanced diversity and similarities about how people from different cultures interpret and assign meaning to their cultural representation. Thus our selection criteria took a purposive sampling approach (Blandford et al., 2016), aiming to identify a diverse set of individuals with various intersecting identities relevant to the local context. We were not prescriptive regarding the exact numbers and form of diversity in our sample, but instead we relied on our partners to guide us towards axes of marginalization and social differentiation that might influence socio-cultural experiences in the country. These led to a sample that was highly contextualized in its diversity but not necessarily exhaustive of all forms of diversity. Minority and diversity here are interpreted in the context of the respective country, for instance, Hinduism is not a minority religion in India, but it is in Sri Lanka and Pakistan, while being Muslim is not a minority in Pakistan, but it is in Sri Lanka. 15 participants per country were selected. Ultimately, 11 participants each from Sri Lanka and Pakistan and 15 from India (spread across two workshops) took part. Participants’ backgrounds are detailed in the appendix.
3.2. Study Activities
@@ 3.2.1. Pre-workshop surveys
We asked each participant to create prompts for AI image generation, drawing on their unique cultural experiences by mixing salient identity markers (e.g., “Punjabi woman”), regional landmark or cultural event, resulting in prompts such as “a group of Punjabi women in front of Wazir Khan Mosque in Lahore.” This ensured both culturally-relevant and diverse prompts, and positioned participants as experts in evaluating the resulting images.
3.2.2. Individual evaluation task
@@
We wanted to give each participant space to individually evaluate the images about their cultures generated by their prompts, before joining the workshop. Thus, after we gathered the prompts from the participants, one of the authors generated images based on each participant’s prompts and compiled a personal collection of prompt and images for each person. Participants were invited to reflect on the image and leave commentary on whether the image was a good representation of their cultural experience. We purposefully did not define “good” or “representation,” since those were the terms we wanted participants to define for us in the workshop.
3.2.3. Workshop
@@
To facilitate collective reflection and discussion on cultural representation in AI images, we convened four-hour workshops across three sites: two in-person workshops in Pakistan and Sri Lanka, and two online workshops in India where logistical constraints necessitated a virtual format. Online workshops were held via Google Meet. Each workshop lasted roughly four hours and had three sections. First, participants engaged in a reflective discussion on their representational goals, drawing from the image-prompt pairs they had evaluated before the workshop. Next, they participated in prompting exercises, iteratively refining prompts through multi-turn interactions with a generative AI model capable of text-to-text, image-to-text, image-to-image, and text-to-image tasks. Finally, participants shared their experiences with image generation and representation in an evaluation discussion. The pre-workshop evaluation task complemented the workshop by providing participants with space for individual reflection, priming them for more informed group discussions. This combination allowed us to explore both individual interpretations and collective understandings of cultural representation, positioning participants as experts in their cultural contexts. Individual evaluations captured participants’ diverse perspectives on AI-generated cultural representation, while the workshops encouraged dialogue to uncover collective interpretations and contestations in meaning-making.
3.3. Data Analysis
@@
All workshops were recorded and transcribed, and where required, translated by authors or local partners. All four authors participated in data preparation (e.g., data cleaning and transcribing) and analysis. We took a reflexive thematic analysis approach to analyze the workshop data, following Braun and Clarke (Braun and Clarke, 2006, 2021)—this includes the transcripts as well as participants’ responses to the individual pre-workshop evaluation task. We met regularly as a group throughout the analysis period to inductively generate themes that captured patterns of shared meaning across workshop sections and workshops (Braun and Clarke, 2021). We used the digital whiteboard Mural to iteratively cluster the codes into larger themes, discussing the relationship between codes and themes as we went and resolving any disagreements in synchronous group discussions. To attribute the quote we used in the following section to the participants, we have used anonymized participant identifiers: country code (IN for India, PK for Pakistan, and LK for Sri Lanka) concatenated with a number.
4. Findings
@@
In section 4.1, we show the granular, multi-dimensional categories of cultural representation people draw on when evaluating images, complicating the singular constructs of representation in thin evaluations. In section 4.2, we explore how participants’ goals for appropriate cultural representation were deeply situated within their specific cultural contexts, dynamically change over time, and were discursively constructed through collective dialogues. In section 4.3, we examine the situated forms of social knowledge that underpinned these evaluations, emphasizing the importance of thick evaluations that draw on people’s lived experiences and cultural expertise.
@@
4.1. People evaluate cultural representation through multi-dimensional categories
@@
When asked to evaluate cultural representativeness of AI-generated images, our participants constructed multiple dimensions of representation to evaluate, demonstrating the need for more fine-grained categories of cultural representation than singular constructs like accuracy of representation (cf. Zhang et al., 2024). We distill these into five dimensions, which point to the richer categories that thick evaluations could evaluate for: incorrectness (the accuracy of depicted physical objects—the closest to existing approaches (e.g., Zhang et al., 2024)), missingness (absence of iconic and expected cultural elements), specificity (whether the subject of the image was specific to their particular sub-culture), coherence (whether all elements of an image were appropriate given cultural and social norms), and connotation (the symbolic meanings and interpretations associated with an image). These expanded categories of cultural representation demonstrate the need for constructs that go beyond the physical world to capture the social worlds of a culture.
4.1.1. Incorrectness
@@
Participants evaluated the (in)correct representation of specific artifacts that have a singular corresponding physical existence outside the image, such as a building or a landmark or a local landscapes. This category has the most congruence with existing thin evaluations, as it seeks to evaluate depictions of artifacts with respect to their existence, and can often be answered with a binary or quantifiable metric since the ‘ground truth’ is more easily determinable than the following dimensions of representation. For instance, participants talked about how the depiction of the Lotus Tower in Sri Lanka can be incorrect or correct since there is only one Lotus Tower with a very particular form. As such, when attempting to generate images, Sri Lankan participants reflected on how they were not able to produce correct images of the Lotus Tower, a famous landmark and the tallest structure in Sri Lanka. However while incorrectness was relevant for evaluating objects with clear counterparts in the physical world, it was not used to assess other forms of cultural representation of the social world, indicating that evaluations of cultural representation encompass more than just factual correctness.
4.1.2. Missingness
@@
Similarly, participants evaluated whether images lacked the expected cultural elements that they felt contributed to a place’s identity—what we refer to as missingness (cf. “erasure:” Qadri et al., 2023a; Shelby et al., 2023; Katzman et al., 2023; Qadri et al., 2025). Participants felt that the absence of common features and structures in images meant to represent their culture meant that a given culture or location’s essence or identity would not be communicated. Thus LK-5, a participant from Sri Lanka, noted scenes that are commonplace in Sri Lanka that were curiously missing across a number of generated images: “iconic places were never represented, like post offices and railway stations. Food items in Sri Lanka were never represented. Beaches were never reflected in any images, [but] we are an island.” In a similar vein, LK-11 pointed out that even nationally significant elements of the Sri Lankan landscape—such as visually distinct species of trees—were absent, but should be represented: “so it should reflect more of our national items if we were to ask it to generate things about Sri Lanka, such as the Naa tree, Bo tree, national bird, animals.” This sense of missingness was closely tied to participants’ ability to distinguish one place from another and capture its unique cultural character—or its iconicity (cf. Barasch and Serrano, 1992).
4.1.3. Specificity
@@
Participants evaluated the specificity of cultural representation—i.e., whether the depiction of cultural elements was specific to their contexts—especially for elements indicative of intersectional identities or subcultures, or cases where cultural artifacts were shared across cultures (but manifested in slightly different ways). For example, when evaluating images of women wearing saris, participants identified the regional nuances of saris that make them specific to their culture: “I would say it’s Marathi. A Marathi sari is a little different. So, the saris also have very different variations. Right. So these saris, particularly makes me feel like it’s from the Northern region. Like that is more, I think, mainstream, India” (IN-3). Evaluating for cultural specificity meant that participants paid close attention to nuanced details that distinguished the cultural practices of their social worlds from others. For other participants, when evaluating cultural representation, it was important to point out that a style of dress was more common in another part of the world than theirs: “this has more Arab cultural influence, like wearing long dresses and wearing hijabs” (PK-8). Similar to missingness, evaluations of specificity were crucial for accurately conveying the unique aspects of a particular culture, bringing with them layers of meaning and symbolism from their social worlds.
4.1.4. Coherence
@@
Participants also evaluated the coherence of the images of their culture—considering the extent to which various elements in an image were in alignment about which culture they were representing and in what ways. They evaluated (in)coherence in multiple ways, such as evaluating images for unrealistic combinations of cultural elements, misaligned behaviors with social and cultural norms, and anachronistic elements of their culture. Participants identified images that contained a mishmash of cultural signifiers that did not typically occur together, such as merging cultural details from different (sub)regions, subcultures, religions, or nationalities into one image: “So the attire is of the people of Tamil. And the lanterns look Chinese, and no symbol of Sinhal[ese] new year” (LK-10). Other participants in this workshop elaborated on the incoherence, pointing out that: “the kids are wearing what normally Sri Lankan kids [are] wearing. But she is also having the [bindhi], and the elder person is kind of Sri Lankan, the mom is kind of European again. And the other kid is having totally different [clothing]—a sari on a male kid I think? So I kind of gave it a two out of five.” (LK-1)
@@
Another form of (in)coherence was temporal, when elements from different time periods were placed together, creating anachronisms. IN-5 noted that some visuals generated in images were correct once, but no longer so: “No one covers their head [now]. My mother used to cover her head. My granny used to cover her head, but I have not seen any woman who’s living in a village or in a rural area [do that] nowadays. It was a part of Punjabi culture, but not anymore.”
4.1.5. Connotations
@@
Finally, participants evaluated the connotations evoked by the images, recognizing that representation in visual media is not just depiction, but also a communicative act. While certain elements of images may be accurate in the sense of potentially occurring in the physical world, participants discussed how those images evoked connotations of the social world that they felt were not appropriate representations of their culture. In the workshop in Pakistan, participants noted that depictions of beards in Pakistani images evoked Western stereotypes about Pakistanis, and they raised concerns about promoting narrow ideas of Pakistani culture: “Everyone has beards in these photos. Many people in Pakistan don’t have a beard. Even in this [workshop] group… this brother does not have a beard, even my beard is so small” (PK-11). Similarly, participants interpreted connotations of poverty from the representation of particular cultural artifacts, which were not inherent to the denotations of the visual images. For example, showing women in saris in India was associated with “a typical traditional thinking about the Indian women since long ancient times.” In some evaluations, participants explicitly expressed frustration as to these images being a signal of ‘them’ seeing ‘us’ in a particular light: “In none of the pictures we see women. I don’t know what they think of us. Like we are from the 19th century and live 200 years ago” (PK-11).
4.2. Goals for cultural representation are situated, dynamic, and negotiated
4.2.1. Goals for representation are situated and dynamic
@@
The previous section showed the more granular categories needed to evaluate the construct of cultural representation in AI images. In this section, we identify how what is considered the goal when evaluating those constructs of cultural representation are developed through a dynamic conversation between the participants about their lived experiences and broader messaging about their culture. Participants emphasized that there was no ’objective’ ideal for representation, but instead what constitutes meaningful cultural representation varies both across and within cultures and contexts. For instance, diversity of representation is one example of a representational goal generative models might aim for, but one participant noted how identities considered diverse in other contexts were, in fact, the dominant identities in her own: “what might be diversity in the First World might actually be monotony in my area… what is diversity in a First World context or in the rest of India is the dominant culture where I come from” (IN-15). For IN-15, when the model attempted to achieve a decontextualized form of diversity from a Western lens, it was inadvertently just depicting dominant cultures: “In the First World, there are attempts at sort of including Muslim representations and all of that… But for example when I give a prompt as a Kashmiri woman, Muslims are the majority. Here [in Kashmir] diversity would look like something else… So here I don’t want just women with hijabs represented. I also want Muslims without hijabs and non-Muslim women represented” (IN-15). Thus, while representational diversity was important for participants, there was no one-size-fits all approach to diversity, as it needed to be contextualized within the social worlds of the users to be meaningful.
@@
4.2.2. Goals for representation are constructed through discursive negotiation
Participants’ goals for and understanding of cultural representation was being actively negotiated in conversation with broader social narratives they had encountered in other visual media and through dialogue with other participants during the workshop. Even seemingly objective judgements of (in)correctness were contextualized by participants within perceived messaging they had encountered outside the specific image at hand. Some participants developed goals for representation in response to stereotypes about their cultures in the media—such as perceptions that South Asian cultures were undeveloped (e.g., PK-11), leading them to want generated images that showed more modernity. Or, when evaluating the failure of models to generate important landmarks from their cultures, participants were contextualizing these failures within perceived general power differentials between Global North and Global South that they had experienced.
@@ 4.2.2. Goals for representation are constructed through discursive negotiation
This tension between goals for representations to combat stereotypes or achieve realism emerged concretely in images that were interpreted as depicting a less modern version of their culture. For instance, in generated images of Pakistani women, some participants noted that women wearing traditional clothes and not more ‘Western clothes’ would convey the impression that Pakistan was not a modern place. However, other participants felt this image still reflected some women, even if it did not represent everyone, and thus was important to retain. For instance, PK-7 argued “ You have to represent the culture as opposed to a small minority of women. To which PK-11 responded, ”But in all pictures [of Pakistani women] we are seeing the same culture. One of the four images could be of a woman from a modern society.” Participants also acknowledged that there may not be consensus on what constituted desired cultural representation, but instead, people’s goals for representation may be based on the communicative intentions of the group being asked. For instance, PK-1 noted “ country officials or ambassadors.. want the best images of Pakistan, they will want [images to] be whitewashed, but we want to have stray cats and [electricity] poles” (PK-1).
@@
4.3. People draw on situated knowledge and experiences to evaluate representation
In this section, we examine the specific types of knowledge participants drew upon for different evaluative categories outlined in Section 4.1. As participants moved from evaluating more empirically demonstrable aspects of the physical world (e.g., is this the Lotus Tower?) to more culturally-situated judgments of the relationship between the physical and social worlds (e.g., is this scene a positive representative of my city?), participants relied on diverse forms of knowledge and deep experience of their social worlds. For instance, evaluations of incorrectness required less cultural and social knowledge, becoming almost a form of pattern recognition. However, other forms of evaluation such as coherence or connotation required more specific knowledge, since they sought to evaluate not just thin concepts of physicality but thicker concepts of sociality. Understanding where social realities were (in)coherently displayed needed an understanding of appropriate behaviors and cultural norms that are often unstated (e.g., what lakes you would boat in and which you would not, or which generation of women might not be covering their head anymore). Even when visual elements of the physical world were being evaluated, they were often linked to social cues relying on hyper-specific visual elements: from the specific type of jewelry you would wear at particular celebrations to recognizing indigenous trees, local hairstyles, how people sit in a particular space, or even the birds that fly above a particular mosque and the potholes that litter a street.
@@ 4.3. People draw on situated knowledge and experiences to evaluate representation
This recognition of outsiders’ limitations also prompted participants to reflect on their own inability to judge the same axes of representation for other cultures. “If someone asks me to evaluate Africa, I already have a bias of Africa and I will evaluate accordingly” (PK-11). While participants emphasized the importance of lived experience for accurate cultural representation, they also recognized the need for this experience to be granular and specific to intra-country social worlds. Even being from the same country, participants admitted that their personal experiences with other subcultures within their own country were often limited. Many shared examples of cultural norms and knowledge that they were not previously familiar with before this workshop. For instance, in the Pakistan workshop, participants discovered different ways of celebrating the holiday of Nauroz, surprised at their own discoveries in conversation with each other:
{quoting}
@@
These findings highlight the deeply situated cultural and social knowledge needed for evaluating the desired or appropriate cultural representation of social worlds in generated images.
5. Discussion

### v2

@@ Methods
In this research, we treat the evaluation process as an object of study itself through community-centered, qualitative methods. These methodological choices respond to increasing calls for broader participation in AI development and design (Delgado et al. 2023; Birhane et al. 2022a). We engaged with participants from three South Asian countries (Sri Lanka, Pakistan, and India) through a series of individual online evaluations and collective in-person workshops. These methods were chosen to allow us to move beyond binary evaluation metrics like thumbs up/down (cf. Collins et al. 2024); we instead sought to capture in-depth qualitative reflections and discussions that participants have as they interpret and evaluate AI-generated images. Given the discursive and collective nature of representation in media, we chose group workshops instead of individual interviews. To preserve participants’ perspectives, we included an individual evaluation task prior to the workshops. Additionally, to ensure the annotation process reflected local cultural norms—including platform choice, framing of questions, and participant composition—we collaborated with local research partners in each region.
Participants and Sites
@@
We chose our research sites for two main reasons. First, we aim to contribute to a growing body of scholarship that seeks to expand the evaluation of AI beyond Western-centric perspectives (Kak 2020; Mohamed et al. 2020; Sambasivan et al. 2021; Qadri et al. 2023; Ghosh et al. 2024b). Second, we intend to move beyond the common “USA vs. India” dichotomy often found in comparative analyses including the Global South (e.g., Raman et al. 2008; Khairullah and Khairullah 2009). To this end, we focused on three South Asian countries: Sri Lanka, Pakistan, and India. They were selected due to the shared cultural histories, while also recognizing the vast diversity within South Asia.
We recruited participants through collaborations with local research partners deeply embedded in their communities. Our partners in Sri Lanka and Pakistan were university researchers with extensive research experience with marginalized local communities, while our partner in India was a research manager experienced in conducting workshops and focus groups. Recruitment began with emails to targeted lists, including local institutions, previous research participants that worked with those partners, and the partners’ professional networks. Interested individuals completed a screening form detailing demographics and background.
@@
Rather than attempting to exhaustively represent any single culture, our aim was to capture nuanced diversity and similarities about how people from different cultures interpret and assign meaning to their cultural representation. Thus our selection criteria took a purposive sampling approach (Blandford et al. 2016), aiming to identify a diverse set of individuals with various intersecting identities relevant to the local context. We were not prescriptive regarding the exact numbers and form of diversity in our sample, but instead we relied on our partners to guide us towards axes of marginalization and social differentiation that might influence socio-cultural experiences in the country. These led to a sample that was highly contextualized in its diversity but not necessarily exhaustive of all forms of diversities. Minority and diversity here are interpreted in the context of the respective country. For instance, Hinduism is not a minority religion in India, but it is in Sri Lanka and Pakistan, while being Muslim is not a minority in Pakistan, but it is in Sri Lanka. 15 participants per country were selected. Ultimately, 11 participants each from Sri Lanka and Pakistan and 15 from India (spread across two workshops) took part. Participants’ backgrounds are detailed in the Appendix.
Study Activities
@@
Pre-workshop surveys. We asked each participant to create prompts for AI image generation, drawing on their unique cultural experiences by mixing salient identity markers (e.g., “Punjabi woman”) with a regional landmark or cultural event, resulting in prompts such as “a group of Punjabi women in front of Wazir Khan Mosque in Lahore.” This fostered diverse and culturally-relevant prompts and enabled participants to be experts in evaluating the resulting images.
@@
Individual evaluation task. We wanted to give each participant space to individually evaluate the images about their cultures generated by their prompts, before joining the workshop. Thus, after we gathered the prompts from the participants, one of the authors generated images based on each participant’s prompts and compiled a personal collection of prompt and images for each person. Participants were invited to reflect on each image and leave commentary on whether the image was a good representation of their cultural experience. We purposefully did not define “good” or “representation,” since those were terms we wanted participants to define themselves during the workshops.
Workshop. To facilitate collective reflection and discussion on cultural representation in AI images, we convened four-hour workshops across three sites: two in-person workshops in Pakistan and Sri Lanka, and two online workshops in India where logistical constraints necessitated a virtual format. Online workshops were held via Google Meet.
@@
Each workshop lasted roughly four hours and had three sections. First, participants engaged in a reflective discussion on their representational goals, drawing from the image-prompt pairs they had evaluated before the workshop. Next, they participated in prompting exercises, iteratively refining prompts through multi-turn interactions with a generative AI model capable of text-to-text, image-to-text, image-to-image, and text-to-image tasks. Finally, participants shared their experiences with image generation and representation in an evaluation discussion.
The pre-workshop evaluation task complemented the workshop by providing participants with space for individual reflection, to prepare for the group discussions. This combination allowed us to explore both individual interpretations and collective understandings of cultural representation, drawing on participants’ role as experts in their cultural contexts. Individual evaluations captured participants’ diverse perspectives on AI-generated cultural representation, while the workshops encouraged dialogue to uncover collective interpretations and contestations in meaning-making.
Data Analysis
@@
All workshops were recorded and transcribed, and where necessary, translated by authors or local partners. All four authors participated in data preparation (e.g., data cleaning and transcribing) and analysis. We took a reflexive thematic analysis approach to analyze the workshop data, following Braun and Clarke (2006, 2021)—this includes the transcripts as well as participants’ responses to the individual pre-workshop evaluation task. We met regularly as a group throughout the analysis period to inductively generate themes that captured patterns of shared meaning across workshop sections and workshops (Braun and Clarke 2021). We used the digital whiteboard Mural to iteratively cluster the codes into larger themes, discussing the relationship between codes and themes as we went and resolving any disagreements in synchronous group discussions. To attribute the quote we used in the following section to the participants, we have used anonymized participant identifiers: country code (IN for India, PK for Pakistan, and LK for Sri Lanka) concatenated with a number.
Findings
@@
We first show the granular, multi-dimensional categories of cultural representation people draw on when evaluating images, complicating the singular constructs of representation in thin evaluations. We then explore how participants’ goals for appropriate cultural representation were deeply situated within their specific cultural contexts, dynamically change over time, and were discursively constructed through collective dialogues. Finally, we examine the situated forms of social knowledge that underpinned these evaluations, emphasizing the importance of thick evaluations that draw on people’s lived experiences and cultural expertise.
@@
People evaluate cultural representation through multi-dimensional categories
@@
When asked to evaluate cultural representativeness of AI-generated images, our participants drew on multiple dimensions of representation to evaluate, demonstrating the need for more fine-grained categories of cultural representation than singular constructs like accuracy of representation. We distill these into five dimensions, which point to the richer categories that thick evaluations could evaluate for: incorrectness (the accuracy of depicted physical objects—the closest to existing approaches (e.g., Zhang et al. 2024)), missingness (absence of iconic and expected cultural elements), specificity (whether the subject of the image was specific to their particular sub-culture), coherence (whether all elements of an image were appropriate given cultural and social norms), and connotation (the symbolic meanings and interpretations associated with an image). These expanded categories of cultural representation demonstrate the need for constructs that go beyond the physical world to capture the social worlds of a culture.
Incorrectness
@@
Participants evaluated the (in)correct representation of specific artifacts that have a singular corresponding physical existence outside the image, such as a building, a landmark, or a local landscape. This category has the most congruence with existing thin evaluations, as it seeks to evaluate depictions of artifacts with respect to their existence, and can often be answered with a binary or quantifiable metric since the ‘ground truth’ is more easily determinable than the following dimensions of representation. For instance, participants talked about how the depiction of the Lotus Tower in Sri Lanka can be incorrect or correct since there is only one Lotus Tower with a very particular form. As such, when attempting to generate images, Sri Lankan participants reflected on how they were not able to produce correct images of the Lotus Tower, a famous landmark and the tallest structure in Sri Lanka. However while incorrectness was relevant for evaluating objects with clear counterparts in the physical world, it was not used to assess other forms of cultural representation of the social world, indicating that evaluations of cultural representation encompass more than just factual correctness.
Missingness
@@
Similarly, participants evaluated whether images lacked the expected cultural elements that they felt contributed to a place’s identity—what we refer to as missingness (see prior work on the related, though distinct, topic of “erasure”: Qadri et al. 2023; Shelby et al. 2023; Katzman et al. 2023; Qadri et al. 2025). Participants felt that the absence of common features and structures in images meant to represent their culture meant that a given culture or location’s essence or identity would not be communicated. LK-5, a participant from Sri Lanka, noted scenes that are commonplace in Sri Lanka that were curiously missing across a number of generated images: “iconic places were never represented, like post offices and railway stations. Food items in Sri Lanka were never represented. Beaches were never reflected in any images, [but] we are an island.” In a similar vein, LK-11 pointed out that even nationally significant elements of the Sri Lankan landscape—such as visually distinct species of trees—were absent, but should be represented: “so it should reflect more of our national items if we were to ask it to generate things about Sri Lanka, such as the Naa tree, Bo tree, national bird, animals.” This sense of missingness was closely tied to participants’ ability to distinguish one place from another and capture its unique cultural character—i.e., its iconicity (cf. Barasch and Serrano 1992).
Specificity
@@
Participants evaluated the specificity of cultural representation—i.e., whether the depiction of cultural elements was specific to their contexts—especially for elements indicative of intersectional identities or subcultures, or cases where cultural artifacts were shared across cultures (but manifested in slightly different ways). For example, when evaluating images of women wearing saris, participants identified the regional nuances of saris that make them specific to their culture: “I would say it’s Marathi. A Marathi sari is a little different. So, the saris also have very different variations. Right. So these saris, particularly, make me feel like it’s from the Northern region. Like that is more, I think, mainstream India” (IN-3). Evaluating for cultural specificity meant that participants paid close attention to nuanced details that distinguished the cultural practices of their social worlds from others. For other participants, when evaluating cultural representation, it was important to point out that a style of dress was more common in another part of the world than theirs: “this has more Arab cultural influence, like wearing long dresses and wearing hijabs” (PK-8). Similar to missingness, evaluations of specificity were crucial for accurately conveying the unique aspects of a particular culture, bringing with them layers of meaning and symbolism from their social worlds.
Coherence
@@
Participants also evaluated the coherence of the images of their culture—i.e., considering the extent to which various elements in an image were in alignment about which culture they were representing and in what ways. They evaluated (in)coherence in multiple ways, such as evaluating images for unrealistic combinations of cultural elements, misaligned behaviors with social and cultural norms, and anachronistic elements of their culture. Participants identified images that contained a mishmash of cultural signifiers that did not typically occur together, such as merging cultural details from different (sub)regions, subcultures, religions, or nationalities into one image: “So the attire is of the people of Tamil. And the lanterns look Chinese, and [there is] no symbol of Sinhal[ese] new year” (LK-10). Other participants in this workshop elaborated on the incoherence, pointing out that: “the kids are wearing what normally Sri Lankan kids [are] wearing. But she is also having the [bindhi], and the elder person is kind of Sri Lankan, the mom is kind of European again. And the other kid is having totally different [clothing]—a sari on a male kid I think? So I kind of gave it a [score of] two out of five.” (LK-1)
@@
Another form of (in)coherence was temporal, when elements from different time periods were placed together, creating anachronisms. IN-5 noted that some elements of images were correct once, but no longer: “No one covers their head [now]. My mother used to cover her head. My granny used to cover her head, but I have not seen any woman who’s living in a village or in a rural area [do that] nowadays. It was a part of Punjabi culture, but not anymore.”
Connotations
@@
Finally, participants evaluated the connotations evoked by the images, recognizing that representation in visual media is not just depiction, but also a communicative act. While certain elements of images may be accurate in the sense of potentially occurring in the physical world, participants discussed how those images evoked connotations of the social world that they felt were not appropriate representations of their culture. In the workshop in Pakistan, participants noted that depictions of beards in Pakistani images evoked Western stereotypes about Pakistanis, and they raised concerns about promoting narrow ideas of Pakistani culture: “Everyone has beards in these photos. Many people in Pakistan don’t have a beard. Even in this [workshop] group… this brother does not have a beard, even my beard is so small” (PK-11). Similarly, participants interpreted connotations of poverty from the representation of particular cultural artifacts, which were not inherent to the denotations of the visual images. For example, showing women in saris in India was associated with “a typical traditional thinking about the Indian women since long ancient times” (IN-4). In some evaluations, participants explicitly expressed frustration as to these images being a signal of ‘them’ seeing ‘us’ in a particular light: “In none of the pictures we see women. I don’t know what they think of us. Like we are from the 19th century and live 200 years ago” (PK-11).
Goals for cultural representation are situated, dynamic, and negotiated
Goals for representation are situated and dynamic.
@@
The previous section showed the more granular categories needed to evaluate the construct of cultural representation in AI images. In this section, we identify how the goals for evaluating those constructs of cultural representation are developed through a dynamic conversation among the participants about their lived experiences and broader messaging about their culture. Participants emphasized that there was no ‘objective’ ideal for representation, but instead what constitutes meaningful cultural representation varies both across and within cultures and contexts. For instance, diversity of representation is one example of a representational goal generative models might aim for, but one participant noted how identities considered diverse in other contexts were, in fact, the dominant identities in her own: “what might be diversity in the First World might actually be monotony in my area… what is diversity in a First World context or in the rest of India is the dominant culture where I come from” (IN-15). For IN-15, when the model attempted to achieve a decontextualized form of diversity from a Western lens, it was inadvertently just depicting dominant cultures: “In the First World, there are attempts at sort of including Muslim representations and all of that… But for example when I give a prompt as a Kashmiri woman, Muslims are the majority. Here [in Kashmir] diversity would look like something else… So here I don’t want just women with hijabs represented. I also want Muslims without hijabs and non-Muslim women represented” (IN-15). Thus, while representational diversity was important for participants, there was no one-size-fits all approach to diversity, as it needed to be contextualized within the social worlds of the users to be meaningful.
@@
Goals for representation are constructed through discursive negotiation.
Participants’ goals for and understanding of cultural representation was actively negotiated in conversation with broader social narratives they had encountered in other visual media and through dialogue with other participants during the workshop. Even seemingly objective judgments of (in)correctness were contextualized by participants within discourses and messages they had encountered outside the specific image at hand. Some participants developed goals for representation in response to stereotypes about their cultures in the media—such as perceptions that South Asian cultures were undeveloped (e.g., PK-11), leading them to want generated images that showed more modernity. Or, when evaluating the failure of models to generate important landmarks from their cultures, participants were contextualizing these failures within perceived general power relations between the Global North and Global South that they had experienced.
@@
This tension between goals for representations to combat stereotypes and realism emerged concretely in images that were interpreted as depicting a less modern version of their culture. For instance, in generated images of Pakistani women, some participants noted that women wearing traditional clothes and not more ‘Western clothes’ would convey the impression that Pakistan was not a modern place. However, other participants felt this image still reflected some women, even if it did not represent everyone, and thus was important to retain. For instance, PK-7 argued “You have to represent the culture as opposed to a small minority of women.” To which PK-11 responded, “But in all pictures [of Pakistani women] we are seeing the same culture. One of the four images could be of a woman from a modern society.” Participants also acknowledged that there may not be consensus on what constituted desired cultural representation, but instead, people’s goals for representation may be based on the communicative intentions of the group being asked. For instance, PK-1 noted that “country officials or ambassadors… want the best images of Pakistan. They will want [images to] be whitewashed, but we want to have stray cats and [electricity] poles” (PK-1).
@@
People draw on situated knowledge and experiences to evaluate representation
In this section, we examine the specific types of knowledge participants drew upon for different evaluative categories outlined in the first section on multi-dimensional categories. As participants moved from evaluating more empirically demonstrable aspects of the physical world (e.g., is this the Lotus Tower?) to more culturally-situated judgments of the relationship between the physical and social worlds (e.g., is this scene a positive representative of my city?), participants relied on diverse forms of knowledge and deep experience of their social worlds. For instance, evaluations of incorrectness required less cultural and social knowledge, becoming almost a form of pattern recognition. However, other forms of evaluation such as coherence or connotation required more specific knowledge, since they sought to evaluate not just thin concepts of physicality but thicker concepts of sociality. Understanding where social realities were (in)coherently displayed needed an understanding of appropriate behaviors and cultural norms that are often unstated (e.g., what lakes you would boat in and which you would not, or which generation of women might not be covering their head anymore). Even when visual elements of the physical world were being evaluated, they were often linked to social cues relying on hyper-specific visual elements: from the specific type of jewelry you would wear at particular celebrations to recognizing indigenous trees, local hairstyles, how people sit in a particular space, or even the birds that fly above a particular mosque and the potholes on a street.
@@
This recognition of outsiders’ limitations also prompted participants to reflect on their own inability to judge the same axes of representation for other cultures. “If someone asks me to evaluate Africa, I already have a bias[ed view] of Africa and I will evaluate accordingly” (PK-11). While participants emphasized the importance of lived experience for accurate cultural representation, they also recognized the need for this experience to be granular and specific to intra-country social worlds. Even being from the same country, participants admitted that their personal experiences with other subcultures within their own country were often limited. Many shared examples of cultural norms and knowledge that they were not previously familiar with before this workshop. For instance, in the Pakistan workshop, participants discovered different ways of celebrating the holiday of Nauroz, surprised at their own discoveries in conversation with each other:
@@
These findings highlight the deeply situated cultural and social knowledge needed for evaluating appropriate cultural representation of social worlds in generated images.
Discussion